# Calibrated and uncertain? Evaluating uncertainty estimates in binary classification models

### v1

@@ Abstract
Rigorous statistical methods, including parameter estimation with accompanying uncertainties, underpin the validity of scientific discovery, especially in the natural sciences. With increasingly complex data models such as deep learning techniques, uncertainty quantification has become exceedingly difficult and a plethora of techniques have been proposed. In this case study, we use the unifying framework of approximate Bayesian inference combined with empirical tests on carefully created synthetic classification datasets to investigate qualitative properties of six different probabilistic machine learning algorithms for class probability and uncertainty estimation: (i) a neural network ensemble, (ii) neural network ensemble with conflictual loss, (iii) evidential deep learning, (iv) a single neural network with Monte Carlo Dropout, (v) Gaussian process classification and (vi) a Dirichlet process mixture model. We check if the algorithms produce uncertainty estimates which reflect commonly desired properties, such as being well calibrated and exhibiting an increase in uncertainty for out-of-distribution data points. Our results indicate that all algorithms are well calibrated, but none of the deep learning based algorithms provide uncertainties that consistently reflect lack of experimental evidence for out-of-distribution data points. We hope our study may serve as a clarifying example for researchers developing new methods of uncertainty estimation for scientific data-driven modeling.
@@ 1 Introduction
In this study we will focus on classification problems. These are tasks where data is sorted into predefined categories based on their features, which are often assumed to be continuous variables. Examples in the natural sciences include determining the gender of fish based on parameters such as length and weight, categorizing patients into positive or negative for a disease based on medical test results or classifying jets produced by the decay of unstable particles in high energy particle physics collision experiments based on jet constituent features. The classifier algorithms we investigate in this study estimate the probability of a new data point belonging to each of the possible classes, as well as the uncertainty of this probability.
@@
The fidelity of the in-distribution estimated class probabilities themselves has also been a topic of much debate. A well calibrated classifier outputs class probability estimates that on average match the empirical class frequency distribution. There is conflicting empirical evidence in the literature on the calibration properties of neural network classifiers. Some studies show that point estimates are not calibrated, raising concerns about their reliability (Guo et al., 2017). Other studies show that neural networks actually are well calibrated (Niculescu-Mizil and Caruana, 2005; Minderer et al., 2021). The field of uncertainty quantification in machine learning is unfortunately fraught with inconsistency, lack of theoretical justification and widespread disagreement on fundamental terms such as probability, uncertainty, and calibration, further exacerbating confusion (Trivedi and Nord, 2025; Perez-Lebel et al., 2023).
The aim of this study is to contribute to a clearer understanding in the field of uncertainty quantification and probabilistic machine learning with a focus on binary classification. We do this by clarifying the concepts of class probability, calibration, uncertainty estimates and out-of-distribution predictions, using the theoretical framework of probability calculus and approximate Bayesian inference. In addition, we have performed a case study where we use these theoretical insights to evaluate qualitative properties of six different probabilistic learning algorithms for class probability and uncertainty estimation: (i) a neural network ensemble (MacKay, 1995), (ii) neural network ensemble with conflictual loss (Fellaji et al., 2024), (iii) evidential deep learning (Sensoy et al., 2018), (iv) neural network with Monte Carlo Dropout (Gal and Ghahramani, 2016), (v) Gaussian process classification (Rasmussen and Williams, 2006) and (vi) a Dirichlet Process Mixture Model (Li et al., 2019). The study investigates the estimates produced by the different algorithms using synthetic data with a clearly defined generating function with smoothly varying densities, inspired by common assumptions made in natural science experiments. Specifically, we address the following three research questions in the case of our data:
@@
The research questions and relevant terminology are explained further in sections 2 and 4. The algorithms tested are described in Appendix A. We hope that our study can serve as an intuitive and easy to implement example of employing toy studies for researchers looking into developing new uncertainty quantification methods using theoretically founded methodology.
@@ 2 Theory
In this section we will describe the theoretical concepts relevant to the case study. We assume the reader is familiar with basic probability theory, Bayesian, and frequentist statistics, but recap the most important aspects before explaining the concept of calibration and out-of-distribution overconfidence. Further we introduce the mean and standard deviation of the posterior distribution as our class probability and uncertainty estimates and explain how these can be calculated.
@@ 2.1 The need for a theory of deep learning
Machine learning techniques such as deep learning, a family of algorithms using artificial neural networks to learn from data and make predictions on unseen data, have proved to be extremely effective tools for problems involving big complex datasets. However, there has traditionally been a lack of interest in establishing mathematical theoretical foundations in the successfully empirically-driven field of deep learning (Breiman, 2001). Despite this, as deep learning has shown itself to be exceptionally effective at solving real-life problems, significant effort has been put into theory development as described in an overview by He and Tao (2020), who classifies attempts at theory development into six different categories. Their category “Stochastic differential equations for modeling stochastic gradient descent and its variants” encompasses the Bayesian approach to the problem, and this is also the approach we use.
@@ 2.2 Probability theory
We consider the following very general model of our data. A dataset with $N$ independent, identically distributed samples $\mathcal{D}_{N}=\{d_{1},d_{2},d_{3},...,d_{N}\}$, with $d_{i}=(\mathbf{c}_{i},\mathbf{x}_{i})$ is created by sampling from the underlying joint probability distribution $p(\mathbf{c},\mathbf{x})$. We let the one-hot vector $\mathbf{c}_{i}$ indicate the binary [^1] Extension to more than two classes is fairly trivial at this point, but turns out to not be as simple in practice, so we will just consider the binary case. class label, a nominal variable, which can take on the values $[1,0]$ (class 1) or $[0,1]$ (class 2). We consider the input features $\mathbf{x}_{i}$ to be continuous $n$-dimensional vectors $\mathbf{x}_{i}=[x_{i}^{1},x_{i}^{2},...x_{i}^{n}]\in\mathbb{R}^{n}$. The aim of the classification problem is to estimate the discrete conditional class probability distribution of a new data point $d_{N+1}$, $\mathbf{P}(\mathbf{c}_{N+1}|\mathbf{x}_{N+1})=[P(c_{N+1}^{1}=1|\mathbf{x}_{N+1}),P(c_{N+1}^{2}=1|\mathbf{x}_{N+1})]$, where $P(c^{1}=1|\mathbf{x})+P(c^{2}=1|\mathbf{x})=1$.
The conditional class probability distribution may be rewritten as a categorical distribution parametrized by $\nu(\mathbf{x})$ as $P(\mathbf{c}|\mathbf{\nu}(\mathbf{x}))=\nu(\mathbf{x})^{c^{1}}\cdot(1-\nu(\mathbf{x}))^{c^{2}}$. We can express $\nu$ in terms of the conditional densities $P(\mathbf{x}|\mathbf{c})$ using Bayes formula
$\nu(\mathbf{x})=P(c^{1}=1|\mathbf{x})=\frac{p(\mathbf{x}|c^{1}=1)p(c^{1}=1)}{p(\mathbf{x}|c^{1}=1)p(c^{1}=1)+p(\mathbf{x}|c^{2}=1)p(c^{2}=1)}.$ (1)
If there is overlap in the conditional densities, classification is not deterministic.
2.3 Calibration properties
The frequency interpretation of probability becomes immediately clear if we consider a dataset sampled from $p(\mathbf{c},\mathbf{x})$ with $N$ samples. The limit of the sample frequencies of each class (the empirical frequency distribution) in any closed set $A_{\epsilon}\in X^{n}$ with volume $\epsilon$ for $\epsilon\rightarrow 0$ and $N\rightarrow\infty$ are equal to the class probabilities $\mathbf{P}(\mathbf{c}|\mathbf{x})$. For this reason we will refer to $\mathbf{P}(\mathbf{c}|\mathbf{x})$ as the long-run frequency distribution (LRFD), although this quantity is often referred to as the true underlying probabilities, true probabilities, real probabilities and the true posterior distribution (Perez-Lebel et al., 2023). The empirical frequency distribution, which is also the maximum likelihood estimate for the Bernoulli distribution, is unbiased as its expectation value for any size dataset equals $\nu(\mathbf{x})$. The calibration literature is focused on evaluating if a classifier’s output matches the empirical frequency distribution of the test set (Nixon et al., 2019; Rubin, 1984). This estimator however, poses a problem when we want to look into the tails of the distribution, as explained in the following.
@@ 2.3 Calibration properties
If we do not have any prior information about the distribution in question we need to choose an uninformative prior which assigns equal probabilities to both classes to do Bayesian inference (Zhu and Lu, 2004; Jaynes, 2003; Bernardo and Smith, 2000). For a Bernoulli distribution with probability of success $\nu$ (observations of class 1), such as in our binary classification task, the conjugate prior is the Beta distribution with density
@@
where $\Gamma$ is the Gamma function and $\alpha,\gamma$ are its parameters. The resulting posterior for an experiment resulting in $N$ trials and $S$ observations of class 1 is also a Beta distribution with density
$p(\nu|N,S)\propto\nu^{\alpha+S-1}(1-\nu)^{\gamma+N-S-1},$
@@
$\mathbb{E}[P(c^{1}=1|\mathbf{x})]=\mathbb{E}[\nu]=\int_{0}^{1}p(\nu|N,S)\nu d\nu=\frac{S+\alpha}{N+\alpha+\gamma}.$
This estimator is consistent as $\lim_{N,S\rightarrow\infty}\mathbb{E}(\nu)=\nu$, but not necessarily unbiased.
To get the desired $\mathbb{E}(\nu)=0.5$ for $N=S=0$ required by an uninformative prior, we have to set $\alpha=\gamma=c$. This gives us a mean and variance of
$\mathbb{E}[P(c^{1}=1|\mathbf{x})]=\mathbb{E}[\nu]=\frac{S+c}{N+2c},$ (2)
$\text{Var}[P(c^{1}=1|\mathbf{x})]=\text{Var}[\nu]=\frac{(N-S+c)(S+c)}{(N+2c+1)(N+2c)^{2}}=\frac{\mathbb{E}[\nu](1-\mathbb{E}[\nu])}{(N+2c+1)}.$ (3)
As explained by Zhu and Lu (2004), the choice of $c$ determines how strong the influence of the prior is on the posterior. Letting $c$ go towards 0 leads to the unbiased maximum likelihood estimator $\lim_{c\rightarrow 0}\mathbb{E}(\nu)=\lim_{c\rightarrow 0}(S+c)/(N+2c)=S/N$, and also the largest possible variance. This prior is essentially two point masses of probability at 1 and 0. If we make one observation of class 1, $S=N=1$, and insert this data into the formula for the posterior, it will (when we approach the limit) shift the expected class frequency completely towards that outcome, which implies a high confidence for the next prediction, contrary to our intuition for an experiment with only one datapoint. Setting $c$ instead to 1 specifies a flat prior over $\nu$ and gives us the Bayes-Laplace estimator $\mathbb{E}(\nu)=(1+S)/(2+N)$, the posterior mean of the probability of class 1 for a new data point will be $\frac{2}{3}$, a much less extreme departure from the $S=N=0$ case. If we want to reduce overconfident estimates for class probability when there is little data, we should choose a prior with stronger bias towards $0.5$. In this case, our estimates will not be well calibrated (unbiased) for few datapoints, but the model will be less prone to high-confidence error. It is therefore necessary to separate the discussion of in-distribution (locally high training statistics) calibration where the bias from the prior is insignificant from the discussion of out-of-distribution (locally low training statistics) predictions where a bias is desirable.

### v2

@@ Abstract
Rigorous statistical methods, including parameter estimation with accompanying uncertainties, underpin the validity of scientific discovery, especially in the natural sciences. With increasingly complex data models such as deep learning techniques, uncertainty quantification has become exceedingly difficult and a plethora of techniques have been proposed. In this case study, we use the unifying framework of approximate Bayesian inference combined with empirical tests on carefully created synthetic classification datasets to investigate qualitative properties of six different probabilistic machine learning algorithms for class probability and uncertainty estimation: (i) a neural network ensemble, (ii) neural network ensemble with conflictual loss, (iii) evidential deep learning, (iv) a single neural network with Monte Carlo Dropout, (v) Gaussian process classification and (vi) a Dirichlet process mixture model. We check if the algorithms produce uncertainty estimates which reflect commonly desired properties, such as being well calibrated and exhibiting an increase in uncertainty for out-of-distribution data points. Our results indicate that all algorithms show reasonably good calibration performance on our synthetic test sets, but none of the deep learning based algorithms provide uncertainties that consistently reflect lack of experimental evidence for out-of-distribution data points. We hope our study may serve as a clarifying example for researchers that are using or developing methods of uncertainty estimation for scientific data-driven modeling and analysis.
@@ 1 Introduction
In this study we will focus on classification problems. These are tasks where data is sorted into predefined categories based on their features, which are often assumed to be continuous variables. Examples in the natural sciences include determining the gender of fish based on parameters such as length and weight, categorizing patients into positive or negative for a disease based on medical test results or event classification in high energy particle physics collision experiments based on the kinematic properties of reconstructed objects. The classifier algorithms we investigate in this study estimate the probability of a new data point belonging to each of the possible classes, as well as the uncertainty of this probability.
@@
The fidelity of the in-distribution estimated class probabilities themselves has also been a topic of much debate. A well calibrated classifier outputs class probability estimates that on average match the empirical class frequency distribution. There is conflicting empirical evidence in the literature on the calibration properties of neural network classifiers. Some studies show that point estimates are not well calibrated, raising concerns about their reliability (Guo et al., 2017). Other studies show that neural networks exhibit reasonably good calibration properties (Niculescu-Mizil and Caruana, 2005; Minderer et al., 2021). The field of uncertainty quantification in machine learning is unfortunately fraught with inconsistency, lack of theoretical justification and widespread disagreement on fundamental terms such as probability, uncertainty, and calibration, further exacerbating confusion (Trivedi and Nord, 2025; Perez-Lebel et al., 2023).
The aim of this study is to contribute to a clearer understanding in the field of uncertainty quantification and probabilistic machine learning with a focus on binary classification. We do this by clarifying the concepts of class probability, calibration, uncertainty estimates and out-of-distribution predictions, using the theoretical framework of probability calculus and approximate Bayesian inference. In addition, we have performed a case study where we use these theoretical insights to evaluate qualitative properties of a representative but limited subset of probabilistic learning algorithms for class probability and uncertainty estimation. Specifically, we evaluated the following algorithms: (i) a neural network ensemble (MacKay, 1995), (ii) neural network ensemble with conflictual loss (Fellaji et al., 2024), (iii) evidential deep learning (Sensoy et al., 2018), (iv) neural network with Monte Carlo Dropout (Gal and Ghahramani, 2016), (v) Gaussian process classification (Rasmussen and Williams, 2006) and (vi) a Dirichlet Process Mixture Model (Li et al., 2019). The study investigates the estimates produced by the different algorithms using synthetic data with a clearly defined generating function with smoothly varying densities, inspired by common assumptions made in natural science experiments. Specifically, we address the following three research questions in the case of our data:
@@
The research questions and relevant terminology are explained further in sections 2 and 4. The algorithms tested are described in Appendix A. We hope that our study can serve as an intuitive and easy to implement example of employing toy studies for researchers looking into developing new uncertainty quantification methods using theoretically founded methodology. We want to make it clear that the study only considers three datasets and a few of the possible uncertainty quantification methods that are available. The results of this study should not be generalized beyond this scope.
@@ 2 Theory
In this section we will describe the theoretical concepts relevant to the case study. We assume the reader is familiar with basic probability theory and both Bayesian and frequentist statistics, but recap the most important aspects before explaining the concept of calibration and out-of-distribution overconfidence. Further we introduce the mean and standard deviation of the posterior distribution as our class probability and uncertainty estimates and explain how these can be calculated.
@@ 2.1 The need for a theory of deep learning
Machine learning techniques such as deep learning, a family of algorithms using artificial neural networks to learn from training data to make predictions on unseen data, have proved to be extremely effective tools for problems involving big complex datasets. However, there has traditionally been a lack of interest in establishing mathematical theoretical foundations in the successfully empirically-driven field of deep learning (Breiman, 2001). Despite this, as deep learning has shown itself to be exceptionally effective at solving real-life problems, significant effort has been put into post-hoc theory development as described in an overview by He and Tao (2020), who classifies attempts at theory development into six different categories. Their category “Stochastic differential equations for modeling stochastic gradient descent and its variants” encompasses the Bayesian approach to the problem, and this is also the approach we use.
@@ 2.2 Probability theory
We consider the following very general model of our data. A dataset $\mathcal{D}_{N}:=\{d_{1},...,d_{N}\}$ with $N$ samples, where $d_{i}=(\mathbf{c}_{i},\mathbf{x}_{i})$, is a finite sample from an infinitely exchangeable sequence [^1] The assumption of exchangeability is slightly weaker than the more common assumption of independent and identically distributed random variables. A sequence is deemed exchangeable if we would assign it the same joint probability regardless of the order of the data points.. We let the one-hot vector $\mathbf{c}_{i}$ indicate a binary [^2] Extension to more than two classes is fairly trivial at this point, but turns out to not be as simple in practice, so we will just consider the binary case. class label, which is considered to be a categorical random measure that can take on the values $[1,0]$ (class 1) or $[0,1]$ (class 2). We consider the input features $\mathbf{x}_{i}$ to be a collection of $d_{x}$ continuous real-valued random measures which can be represented by the $d_{x}$-dimensional vector $\mathbf{x}_{i}=[x_{i}^{1},x_{i}^{2},...x_{i}^{d_{x}}]\in\mathbb{R}^{d_{x}}$. The aim of the classification problem is to calculate the predictive posterior conditional class probability distribution for a new data point, $d_{N+1}=(\mathbf{c},\mathbf{x})$}. This probability distribution is a binary discrete distribution given by $\mathbf{P}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N})=[P(c^{1}=1|\mathbf{x},\mathcal{D}_{N}),P(c^{2}=1|\mathbf{x},\mathcal{D}_{N})]$, where ${P(c^{2}=1|\mathbf{x},\mathcal{D}_{N})}=1-{P(c^{1}=1|\mathbf{x},\mathcal{D}_{N})}$.
By using de Finetti’s representation theorem, it can be shown that if $\{\mathbf{c}_{1},...,\mathbf{c}_{N}\}$ is a finite sample from an infinitely exchangeable sequence of measurements of the binary random quantity $\mathbf{c}_{i}\in\{[1,0],[0,1]\}$, then the joint probability of this data is given by the marginalization over the parameterized distribution
$\begin{split}P(\mathbf{c}_{1},...,\mathbf{c}_{N})=\int_{0}^{1}P(\mathbf{c}_{1},...,\mathbf{c}_{N}|\nu_{\infty}=\nu)p(\nu_{\infty}=\nu)d\nu=\\ \int_{0}^{1}\prod_{i=1}^{N}{\nu}^{c^{1}_{i}}(1-\nu)^{c^{2}_{i}}p(\nu)d\nu,\end{split}$ (1)
where $0\leq\nu\leq 1$. The likelihood $P(\mathbf{c}_{1},...,\mathbf{c}_{N}|\nu_{\infty}=\nu):=P(\mathbf{c}_{1},...,\mathbf{c}_{N}|\nu)$ can be interpreted as the joint distribution of a set of independent measurements that are assumed sampled from a Bernoulli distribution with parameter $\nu$. The probability density distribution ${p(\nu_{\infty}=\nu):}=p(\nu)$ is the prior for the long run frequency distribution (LRFD) of the data. For binary data the LRFD is completely specified by the scalar parameter $\nu_{\infty}=\lim_{N\rightarrow\infty}\sum_{i=1}^{N}c^{1}_{i}/N$. The derivations can be found in Bayesian statistics textbooks such as Bernardo and Smith (2000).
Assuming that the conditional class probability can be defined for every $\mathbf{x}\in\mathbb{R}^{d_{x}}$, we can then express the predictive posterior conditional class probability of class 1 as
$\begin{split}P(c^{1}=1|\mathbf{x},\mathcal{D}_{N})=\int_{0}^{1}P(c^{1}=1|\nu(\mathbf{x}))p(\boldsymbol{\nu}(\mathbf{x})|\mathcal{D}_{N})d\nu(\mathbf{x})=\int_{0}^{1}\nu(\mathbf{x})p(\boldsymbol{\nu}(\mathbf{x})|\mathcal{D}_{N})d\nu(\mathbf{x})\\ =\int_{0}^{1}\nu(\mathbf{x})\frac{p(\mathcal{D}_{N}|\nu(\mathbf{x}))p(\nu(\mathbf{x}))}{p(\mathcal{D}_{N})}d\nu(\mathbf{x})=\mathbb{E}_{\nu\sim p(\nu(\mathbf{x})|\mathcal{D}_{N})}[\nu(\mathbf{x})].\end{split}$ (2)
where $0\leq\nu(\mathbf{x})\leq 1$. $P(c^{2}=1|\mathbf{x},\mathcal{D}_{N})$ is simply $1-P(c^{1}=1|\mathbf{x},\mathcal{D}_{N})$.
In Bayesian statistics, the class probability is a measure of how plausible we would think it is that the data point belongs to that class. Probabilities are always conditioned on all information relevant to the problem, which may or may not include a dataset. In binary classification, $P(c^{1}=1|\mathbf{x})=1$ means we are completely certain that the data point belongs to class 1. An assigned probability of 0.5 means we consider it to be equally likely that it belongs to class 1 or class 2, so we are very uncertain about the classification. In frequentist/classical statistics the class probability is, by definition, the LRFD and we must estimate it empirically. The LRFD of classification problems is often referred to as the true underlying probabilities, true probabilities, real probabilities and the true posterior distribution in machine learning literature (Perez-Lebel et al., 2023). The situation gets more complicated when the class probability is taken to mean the LRFD of another process than the one actually generating the data, such as that of an “unbiased” dataset, where the bias is unknown. We will use the epistemic Bayesian definition of probability as a measure of plausibility conditioned on the data and information we actually have, while also considering the viewpoint of the more commonly used frequentist methodology.
2.3 Calibrated classifiers
Calibration is an important frequentist concept in the uncertainty quantification literature found in machine learning research. We define a perfectly calibrated classification algorithm to be one which outputs estimated class probabilities, $P(c^{1}=1|\mathbf{x},\mathcal{D}_{N})$, that are equal to the LRFD, $\nu_{\infty}(\mathbf{x})$, as defined above. It is by definition impossible to empirically validate that a classifier has this property. The calibration literature is therefore focused on instead evaluating if a classifier’s output matches the empirical class frequency distribution, $\hat{F}_{N}=[\hat{\nu}_{N},1-\hat{\nu}_{N}],\hat{\nu}_{N}=\sum_{i=1}^{N}\frac{c^{1}_{i}}{N}$, of the test set (Nixon et al., 2019; Rubin, 1984).
This approach can be justified mathematically as the empirical class frequency distribution is the unbiased maximum likelihood estimator for a dataset consisting of random binary variables assumed sampled from a Bernoulli distribution. In our case, unbiased means that the expected value of the class probability estimate, with the expectation taken over all possible variations of a finite dataset, is equal to the LRFD. This can be easily derived
$\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}|\nu_{\infty})}\big[\sum_{i=1}^{N}\frac{c^{1}_{i}}{N}\big]=N\sum^{c^{1}\in{0,1}}\frac{c^{1}}{N}P(c^{1}|\nu_{\infty})=1\times\nu_{\infty}+0\times(\nu_{\infty}-1)=\nu_{\infty}.$ (3)
As we are interested in conditional class probabilities, and our test dataset is finite, we have to evaluate calibration over bins in $(P,\mathbf{x})$-space. A classifier can be judged to be locally “well calibrated” if the sample mean of the estimated class probabilities in a specific bin are not too far from the empirical class frequency distribution calculated over the same bin.
Common tools for evaluating calibration are reliability diagrams which do the binning over $P$, and global metrics such as the expected calibration error (ECE)(Guo et al., 2017) which measure the average calibration over all the bins in the reliability diagram. These may pick up on miscalibration if the data points in each bin are clustered together. Global metrics for calibration such as approximations of the Kullback-Leibler divergence and the Wasserstein distance, evaluated over the entire test set, will be dominated by in-distribution data points and are therefore not sensitive to outliers in the data.
2.4 Asymptotic properties of calibrated classifiers
Using calibration metrics, through $\hat{\nu}_{N}$, to validate the predictive posterior conditional class probabilities is equivalent to assuming that we are calculating the probabilities in Equation 2 using a specific prior called the Haldane’s prior. For cases where we have a lot of data, this is an unproblematic choice which is also useful from the perspective of Bayesian statistics as we want our models to converge towards the LRFD (Rubin, 1984). However in the cases where we only have a few data points, or none, we show that the theoretical estimates we get by using this prior does not align with the stated desired properties of uncertainty quantification.
@@ 2.4 Asymptotic properties of calibrated classifiers
If we do not know anything about how the classes are distributed, we should choose an uninformative prior, $p(\nu)$, that assigns equal probabilities to both classes (Zhu and Lu, 2004; Jaynes, 2003; Bernardo and Smith, 2000). For a Bernoulli distribution the conjugate prior is the Beta distribution with density
@@
where $\Gamma$ is the Gamma function and $\alpha,\gamma$ are its parameters. The resulting posterior for an experiment resulting in $N$ trials and $s_{N}$ observations of class 1 is also a Beta distribution with density
$p(\nu|\mathcal{D}_{N})=p(\nu|N,s_{N})\propto\nu^{\alpha+s_{N}-1}(1-\nu)^{\gamma+N-s_{N}-1},$
@@
$P(c^{1}=1|\mathcal{D}_{N})=\mathbb{E}_{\nu\sim p(\nu|\mathcal{D}_{N})}[\nu]=\int_{0}^{1}\nu p(\nu|N,s_{N})d\nu=\frac{s_{N}+\alpha}{N+\alpha+\gamma}.$
This estimator converges to $\nu_{\infty}$ in the limit of infinite data, but is only unbiased if $\alpha=\gamma=0$.
To get the desired equal class probabilities for $N=s_{N}=0$, we have to set $\alpha=\gamma$. This gives us a mean and variance of
$P(c^{1}=1|\mathcal{D}_{N}))=\mathbb{E}_{\nu\sim p(\nu|\mathcal{D}_{N})}[\nu]=\frac{s_{N}+\alpha}{N+2\alpha},$ (4)
$\text{Var}_{\nu\sim p(\nu|\mathcal{D}_{N})}[\nu]=\frac{(N-s_{N}+\alpha)(s_{N}+\alpha)}{(N+2\alpha+1)(N+2\alpha)^{2}}=\frac{\mathbb{E}[\nu](1-\mathbb{E}[\nu])}{(N+2\alpha+1)},$ (5)
where we will from now on use the shorthand notation $\mathbb{E}[\nu]$ to mean the expectation taken over the posterior distribution for $\nu$ and not over possible variations of the dataset.
As explained by Zhu and Lu (2004), the choice of $\alpha$ determines how strong the influence of the prior is on the posterior. Letting $\alpha$ go towards 0 leads to the unbiased maximum likelihood estimator $\lim_{\alpha\rightarrow 0}\mathbb{E}(\nu)=\lim_{\alpha\rightarrow 0}(s_{N}+\alpha)/(N+2\alpha)=s_{N}/N$. $\text{Beta}(0,0)$ is the Haldane prior and is an improper prior with all the probability mass concentrated in two point masses at 0 and 1. If we make one observation of class 1, $s_{N}=N=1$, and insert this data into the formula for the posterior, the resulting class probability and variance for class 1 are 1 and 0 respectively. This implies the highest possible confidence that the next datapoint will also be of class 1, contrary to our intuition for an experiment with only one datapoint.
Setting $\alpha$ to 1 specifies a flat prior over $\nu$ and gives us the Bayes-Laplace estimator $\mathbb{E}(\nu)=(1+s_{N})/(2+N)$. The posterior mean of the probability of class 1 for a new data point will be $\frac{2}{3}$, a much less extreme departure from 0.5. If we want to reduce overconfident estimates for class probability when there is little data, we should set a prior with stronger bias towards $0.5$ by choosing a larger value for $\alpha$ in the prior. If we do this, our estimates will by definition never be perfectly calibrated (unbiased in the frequentist/classical statistics sense), but the model will be less prone to high-confidence error. It is therefore necessary to separate the discussion of in-distribution (locally high training statistics) calibration, where the bias from the prior is insignificant, from the discussion of out-of-distribution (locally low training statistics) predictions, where a bias is desirable.